### v1

# Prompt-SAW: Leveraging Relation-Aware Graphs for Textual Prompt Compression

Abstract
Large language models (LLMs) have shown exceptional abilities for multiple different natural language processing tasks. While prompting is a crucial tool for LLM inference, we observe that there is a significant cost associated with exceedingly lengthy prompts. Existing attempts to compress lengthy prompts lead to sub-standard results in terms of readability/interpretability of the compressed prompt, with a detrimental impact on prompt utility. To address this, we propose Prompt-SAW: Prompt compresSion via Relation AWare graphs, an effective strategy for prompt compression over task-agnostic and task-aware prompts. Prompt-SAW uses the prompt’s textual information to build a graph, later extracts key information elements in the graph to come up with the compressed prompt. We also propose GSM8k-aug, i.e., an extended version of the existing GSM8k benchmark for task-agnostic prompts in order to provide a comprehensive evaluation platform. Experimental evaluation using benchmark datasets shows that prompts compressed by Prompt-SAW are not only better in terms of readability, but they also outperform the best-performing baseline models by up to 14.3% and 13.7% respectively for task-aware and task-agnostic settings while compressing the original prompt text by 33.0% and 56.7%.
1 Introduction
Large Language Models (LLMs) have attracted considerable attention for their superior performance across a wide range of applications. For this, instructions (aka. prompts) play a crucial role in extending the capabilities of LLMs for multiple different tasks. The prompts provide the provision to guide the model to elucidate desired model behavior without perturbing the model parameters. This is also highlighted in recent studies that show well-designed prompts and the integration of external knowledge are significant to enhance the effectiveness of LLMs’ (Sahoo et al., 2024). Different LLMs-related techniques directly benefiting from prompts include but are not limited to: In-Context Learning (Dong et al., 2022), Chain-of-Thought (Wei et al., 2022), Retrieval Augmented Generation (Lewis et al., 2020), and Agents (Park et al., 2023) etc. Generally, prompts may be sub-divided into two types: task-aware and task-agnostic prompts, a quick overview is given in Appendix A.2 and Appendix A.3 respectively.
At the same time, the abilities of LLMs are significantly compromised/constrained by increasingly lengthy prompts, even comprising thousands of tokens. Lengthy prompts not only obscure requisite information but also increase computational costs and incur inference latency. To tackle this challenge, prompt compression techniques, e.g., (Li, 2023), have garnered significant interest. These approaches are based on the fact that natural language is inherently redundant (Shannon, 1951). Thus, it is possible to substantially compress the length of original textual prompts by preserving requisite information in small segments.
Existing prompt compression approaches focus on compressing text at the token level, i.e., they verify whether compression is applicable to each individual token. For instance, Li (2023) proposed Selective-Context that uses a compact language model to evaluate context’s lexical units, enabling compression by eliminating units with minimal information. Also, LLMLingua (Jiang et al., 2023a) and LongLLMLingua (Jiang et al., 2023b) developed budget control mechanisms to compresses prompts based on their perplexity.
While existing approaches could enhance the ability to deal with lengthy prompts for LLMs, they lack grammatical coherence, i.e., existing approaches neglect the syntactic and semantic structure of the compressed prompt. This is because contemporary prompt compression methods primarily focus on quantifying token-level information, neglecting the overall grammatical structure of the compressed prompt. Such ignorance not only increases the risk of semantic loss within the compressed prompt but also hampers its readability for human readers. An example in this regard is shown in Figure 1, where the original prompt text: ”Two women have won the prize: Curie and Maria Goeppert-Mayer” is compressed to: ”won twoes forg01 theate women prize:ertMayer” by LongLLMlingua Jiang et al. (2023b).
Figure 1: An exmple of prompt compression example of previous token-level based method LongLLMlingua (Jiang et al., 2023b)
To fill in the gap, in this paper, we propose Prompt-SAW, i.e., Prompt compresSion via Relation AWare graphs, a novel method designed to cut down unnecessary information in the prompt text by using Knowledge Graph (KG) structures to exploit the small-scale information elements (Section 3.1) in the prompts, i.e., information units comprising entities and their underlying relations.
Prompt-SAW first extracts all entities and their relations in the prompt to formulate the graph. Later, (i) for task-aware prompts, Prompt-SAW looks for small-scale information elements in the graph to only retain task-specific information as a sub-graph, (ii) for task-agnostic prompts, Prompt-SAW measures similarity scores between successive information elements in the graph to remove the redundant elements to obtain required sub-graph. To retain the syntactic and semantics of the prompt structure, Prompt-SAW finally reinstates the information contained in the sub-graph resulting in an optimized and compressed prompt.
We conducted extensive experimental analysis of Prompt-SAW under both task-agnostic and task-aware settings against existing best-performing models as baselines. For evaluation, we used: (i)  GSM8k-aug, i.e., an extended experimental setting proposed by us for open-source benchmark GSM8K (Cobbe et al., 2021), and (ii) open-source benchmark NaturalQuestions (Liu et al., 2023). Experimental results show that Prompt-SAW significantly outperforms other baseline models. We summarize the key contributions of this work as follows:
  • We propose Prompt-SAW, a novel framework crafted for compressing prompts by exploiting graph structures to infer key information elements in the prompt that are helpful for compression.
  • As current benchmarks for task-agnostic prompts lack comprehensive evaluation, we propose GSM8k-aug, an extended version of existing GSM8k benchmark for an intensive evaluation of Prompt-SAW.
  • We demonstrated the effectiveness of Prompt-SAW by comprehensive experiments showing Prompt-SAW attained state-of-the-art performance outperforming baseline models by up to 14.3% and 13.7% respectively for task-aware and task-agnostic settings while compressing the original prompt text by 33.0% and 56.7%.
2 Related Work
Prompt Compression. Prompt compression techniques are used to reduce the inference cost of LLMs across a wide range of applications. Existing work can be categorized into soft prompt compression and discrete prompt compression.
Soft prompts were introduced by Lester et al. (2021). A soft prompt integrates additional trainable parameters at the model’s input stage. (Wingate et al., 2022) emphasized that soft prompt compression effectively retains crucial abstract information with a reduced parameter count. Xu et al. (2023b) emphasized that carefully crafted prompts are helpful in augmenting the end-performance of compressed LLMs, also the compressed LLMs are helpful in the prompt learning phase.
Compared to soft prompt compression, discrete prompt compression seeks to optimize the effectiveness of prompts via token-level search strategies. Jung & Kim (2023) employed policy networks to eliminate unnecessary tokens for prompt compression. Li (2023) utilized self-information metrics to identify and remove superfluous information in prompts. Capitalizing on these advancements, Jiang et al. (2023a) and Jiang et al. (2023b) formulated algorithms for dynamically adjusting compression rates across different prompt sections, giving precedence to tokens with higher perplexity.
Despite the significant advancements achieved by these studies, their primary focus lies on token-level compression, neglecting the comprehensive graph structure information inherent in the prompt.
Knowledge Graphs (KGs) for LLM. KGs organize information as structured units, i.e., relational triplets (explained in Appendix A.1), that encapsulate a wide variety of entities/concepts along with underlying relations (Ji et al., 2020). Pan et al. (2023) illustrated multiple different scenarios for integration of KGs with LLM for knowledge and data-driven bi-directional reasoning. Luo et al. (2023) combined LLMs with KGs for interpretable reasoning over KG Question Answering tasks. Kim et al. (2023) introduced an innovative framework that leverages LLM’s reasoning capabilities for executing KG-based tasks. To the best of our knowledge, Prompt-SAW is the first to make an attempt to leverage knowledge graph structure for prompt compression.
3 Preliminaries
In this section, we first introduce mathematical notations and formulate our problem. Background on the core concepts required for the design and development of Prompt-SAW is provided in Appendix A.
3.1 Notations.
We use $P$ and $C$ to represent the original and compressed prompt respectively. Likewise, we use $N$ and $\widetilde{N}$ to represent the length of the original and compressed prompt. We use $\eta=\widetilde{N}/N$ to represent the compression rate and $1/\eta$ as the compression ratio. $\eta^{*}$ is used to represent the target compression rate. The graph is represented by $\mathcal{G}=\{(e_{i},r_{i},e_{i}^{{}^{\prime}})\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$, where $e_{i}$, $r_{i}$ and $e_{i}^{{}^{\prime}}$ represent the subject entity, relation and object entity in the graph respectively; =$\mathcal{E}=\{e_{1},e_{2},\cdots,e_{m}\}$ and =$\mathcal{R}=\{r_{1},r_{2},\cdots,r_{n}\}$ denote the set of entities and relations in $\mathcal{G}$. $g_{i}=(e_{i},r_{i},e_{i}^{{}^{\prime}})$ is used to represent small-scale information elements in $\mathcal{G}$, this is equivalent to graph triplet. $M$ represents auxiliary models used for graph construction. $E$ is used to represent the encoder network. $\delta$ represents the similarity threshold used for sub-graph construction.
3.2 Problem Setting
In this work, we aim to design and develop an effective prompt compression strategy that can cut down the prompt text by only preserving the requisite information content while at the same time maintaining the semantics and end performance of the prompt to the best possible extent.
Formally, we aim to generate a compressed prompt $C=\{c_{i}\}_{i=1}^{\widetilde{N}}$ given the original prompt $P=(p^{\mathrm{ins}},p^{\mathrm{info}},p^{\mathrm{que}})$, where $p^{\mathrm{ins}}=\{p_{i}^{\mathrm{ins}}\}_{i=1}^{N^{\mathrm{ins}}}$, $p^{\mathrm{info}}=\{p_{i}^{\mathrm{info}}\}_{i=1}^{N^{\mathrm{info}}}$, and $p^{\mathrm{que}}=\{p_{i}^{\mathrm{que}}\}_{i=1}^{N^{\mathrm{que}}}$, denote the prompt instruction, information and question, respectively; $\widetilde{N}$, $N^{\mathrm{ins}}$, $N^{\mathrm{info}}$ and $N^{\mathrm{que}}$ represent the number of tokens in $C$, $p^{\mathrm{ins}}$, $p^{\mathrm{info}}$, $p^{\mathrm{que}}$ and respectively. We denote $N=N^{\mathrm{ins}}$ + $N^{\mathrm{info}}$ + $N^{\mathrm{que}}$ as the length of the original prompt.
4 Prompt-SAW
In this section, we provide details of Prompt-SAW. The workflow is shown in Figure 2. As shown in the figure, Prompt-SAW takes the original prompt text as input and generates the compressed prompt as the output.
In contrast to the existing token-level compression methods, in Prompt-SAW we use a graph structure to effectively represent the textual information in the prompt, which is helpful to analyze the key aspects of the prompt. Later, we can refine the information in the graph structure to come up with a compressed prompt in a way that: (i) The semantic consistency of the compressed prompt is preserved; (ii) The end performance and/or utility of the prompt is not distorted. Below, we first introduce the motivation of Prompt-SAW, followed by the prompt compression process.
4.1 Motivation of Prompt-SAW
Prompt-SAW is motivated by the observation that the key information within the prompt text could be inferred as a set of entities and relations, which can also be organized into a graph structure, commonly known as a knowledge graph in literature.
Formally, given a prompt text $P$, we claim it encompasses a set of entities $(\mathcal{E})$, i.e., names of persons, locations, organizations, miscellaneous elements, etc., (Ali et al., 2020; 2021). These entities serve as the key elements of the prompt structure. In addition to the entities, we can also infer some relations $\mathcal{R}$ in $P$ that may be used to describe the connections between the entities. Prompt-SAW re-organizes these key elements of the prompt (i.e., entities and their relations) in a graph structure, represented by $\mathcal{G}=\{(e_{i},r_{i},e_{i}^{{}^{\prime}})\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$. We use $g_{i}=(e_{i},r_{i},e_{i}^{{}^{\prime}})$ to represent the $i$-th information element of $\mathcal{G}$, i.e., a fact stating that $e_{i}$ has $r_{i}$-$th$ relation with $e_{i}^{{}^{\prime}}$.
We argue this transformation of text information to graph is a more reasonable and natural approach as: (i) It helps in highlighting the key information elements in the prompt. (ii) Later, analyzing these key entities in combination with underlying relations helps in filtering/digging out the salient content within the prompt to come up with a compressed prompt.
Figure 2: An illustration of Prompt-SAW.
4.2 Workflow of Prompt-SAW
The workflow of Prompt-SAW consists of two parts. First, it uses the information in prompt $P$ to construct a graph $\mathcal{G}$. Then, based on the specific scenario:
(a) Task-aware scenario. For this scenario, we traverse the graph $(\mathcal{G})$ in a way to preserve only the information elements that are relevant to the task as task-specific subgraphs, indicative of information useful for the compressed prompt.
Algorithm 1 Subgraph Extraction
1:
2: #$\eta^{*}:$ Target compression rate
[…]
4: #$len():$ Compute the length of graph structure, as the sum of individual tokens.
5:subgraph $\mathcal{G}_{subset}$
6:$\mathcal{G}_{subset}=\{\}$
7:for $i\in Index_{\text{ranked}}$ do
8:     $\mathcal{G}_{subset}.insert(g_{i})$
9:      #compute compression rate
10:     $Rate=len(\mathcal{G}_{subset})/len(\mathcal{G})$
11:      #break if meet the constraint
12:     $\mathbf{If}$ $Rate>\eta^{*}$ then
13:     $\qquad Break$
14:     end If
15:end for
16:return $\mathcal{G}_{subset}$
Algorithm 2 Binary Search
1:
2: #$\eta^{*}:$ Target compression rate
[…]
4: #$\gamma:$ interval threshold
5:subgraph $\mathcal{G}_{subset}$
6:double $l=0,r=1$
7:while $r-l>\gamma$ do
8:     double $mid=(l+r)/2$
9:     $\mathcal{G}_{subset}=Compress(\mathcal{G},mid)$
10:     $Rate=len(\mathcal{G}_{subset})/len(\mathcal{G})$
11:     $\mathbf{If}$ $Rate>\eta^{*}$ then
12:     $\qquad r=mid$
13:     Else
14:     $\qquad l=mid$
15:end while
16:$\delta=(l+r)/2$ compression threshold
17:$\mathcal{G}_{subset}=Compress(\mathcal{G},\delta)$
18:return $\mathcal{G}_{subset}$
(b) Task-agnostic scenario. For this scenario, we no longer have access to task-specific information. Thus, we use similarity scores between the information elements in $\mathcal{G}$ to identify and remove the redundant elements to obtain subgraphs that are helpful for compression.
We will provide details in the following subsections.
4.2.1 Graph Construction.
For graph construction from the text data, we use an in-context learning prompt that prompts a language model to construct the graph from the original prompt text as follows.
$\mathcal{G}={M}(P_{\text{template}}(P)),$ (1)
where $(P)$ is the prompt text and $P_{\text{template}}$ is the prompt template (explained in Appendix B.1) used to guide the model (M) to extract the graph structure $(\mathcal{G})$. For graph construction, we typically prefer a small-scale model for $M$, as a lesser number of parameters can avoid a higher computational cost incurred by large models, which eventually contributes to the overall compression time overhead. It is notable that in addition to the above-mentioned method, we also employed traditional approaches for graph construction, e.g., OpenIE (Pal & Mausam, 2016), however, these approaches yield relatively lower performance.
4.2.2 Task-aware Prompts
Task-aware scenarios refer to the settings when the information within the prompt is helpful and/or is related to the end-task, e.g., question answering. For such cases, Prompt-SAW aims to retain only the task-specific information in $\mathcal{G}$, while filtering out the redundant/useless information. For this, it first uses an encoder function to get the embeddings for the prompt question, as follows.
$Emb_{p^{\mathrm{que}}}=E(p^{\mathrm{que}})$ (2)
where $Emb_{p^{\mathrm{que}}}$ is the embedding for prompt question $(p^{\mathrm{que}})$, and $E$ is the encoder network. Then, it computes the pair-wise similarity between the $Emb_{p^{\mathrm{que}}}$ and information elements in $\mathcal{G}$, as shown below.
$Sim_{\mathcal{G}}=\{E(g_{i})\cdot Emb_{p^{\mathrm{que}}}\;|\forall\;g_{i}\in\mathcal{G}\}$ (3)
where $g_{i}$ corresponds to the $i$-$th$ information element in $\mathcal{G}$, $E(g_{i})$ is used to encode the information in $g_{i}$, $Sim_{\mathcal{G}}$ is the set of the similarity scores between information element in $\mathcal{G}$ and the question embeddings $Emb_{p^{\mathrm{que}}}$. Later, it ranks the scores in $Sim_{\mathcal{G}}$ in order to retain only the elements in $\mathcal{G}$ showing a higher degree of similarity with $p^{\mathrm{que}}$, as shown below.
$Index_{\text{ranked}}=\text{Rank}(Sim_{\mathcal{G}})$ (4)
where $\text{Rank}(\cdot)$ is used to sort the similarity scores in $Sim_{\mathcal{G}}$ and return corresponding high-ranked information elements as $Index_{\text{ranked}}$. We then use the information in $Index_{\text{ranked}}$ to iterate over $\mathcal{G}$ to extract the sub-graph $\mathcal{G}_{\text{subset}}$ not surpassing the targeted compression ratio $\eta^{*}$. Its process-flow is illustrated in Algorithm 1.
Workflow of Algorithm 1. The workflow of Algorithm 1 is explained as follows: (i) initialize $\mathcal{G}_{subset}$ as an empty set (line-1); (ii) for each element in $Index_{\text{ranked}}$ repeatedly add $g_{i}$ in $\mathcal{G}_{subset}$ until the compression rate surpasses the target compression rate {*} (lines 2-7); (iii) return final graph $\mathcal{G}_{subset}$ as output (line-10).
Finally, we restore/reconstruct the information elements in $\mathcal{G}_{\text{subset}}$ to come up with our compressed prompt $C$, as shown below.
$C=e_{1}\oplus r_{1}\oplus e_{1}^{\prime}\oplus\text{[SEP]}\oplus\cdots e_{i}\oplus r_{i}\oplus e_{i}^{\prime}\oplus\text{[SEP]}\oplus\cdots e_{n}\oplus r_{n}\oplus e_{n}^{\prime}$ (5)
where $\oplus$ is the concatenation operator used to combine the entities and relations within the information elements $(g_{i})$ in the extracted subgraph $\mathcal{G}_{\text{subset}}$, and [SEP] is the delimiter used to separate different information elements in $\mathcal{G}_{\text{subset}}$.
4.2.3 Task-agnostic Prompts
A task-agnostic scenario implies that it is almost impossible to filter useful and/or task-specific information within the original prompt text $(P)$. In such cases, Prompt-SAW looks for recurring information elements in $P$ for probable prompt compression. We assume two main sources of recurring elements in $P$, i.e., (i) the verbose expression of the prompt itself and (ii) the repeated element generated by auxiliary models. Note that these assumptions are based on empirical observation illustrating that large models’ re-reading phenomenon leads to the repeated generation of the extracted knowledge (Yan et al., 2023).
For compression over task-agnostic scenarios, we sequentially traverse the information elements in $\mathcal{G}$ and select only the elements exhibiting a lower similarity with priorly selected information elements. Our underlying intuition is that highly similar information elements will carry repeated information. Thus, we could avoid redundant information in $P$ by selecting only dissimilar elements.
For this, we use a threshold $\delta$ as a selection criteria for Prompt-SAW. The value of the $\delta$ is determined using a binary search algorithm (shown in Algorithm 2) that computes an appropriate value of threshold $\delta$ required to meet the targeted compression rate $\eta^{*}$.
Workflow of Algorithm 2. The process-flow of Algorithm 2 is explained as follows: (i) firstly, we initialize an interval $[l,r]$ for the threshold $\delta$ (line-1); (ii) at each step, we partition the interval into two parts $[l,mid]$ and $[mid,r]$ via the midpoint $mid=(l+r)/2$ (line-3); (iii) we will compute the graph subset, i.e., $\mathcal{G}_{subset}$ via function $Compress()$ (explained below) with the value of $mid$ as threshold, shown in line-4; (iv) compute the compression rate for the $\mathcal{G}_{subset}$ and accordingly update the values of $l$ and $r$ (lines 6-9). Specifically, if the compression rate is smaller than $\eta^{*}$, then the current threshold is too stringent thus we judge $\delta$ is in $[mid,r]$, otherwise it is in $[l,mid]$; (v) depending upon the interval threshold $\gamma$ (line-2), we compute the value $(l+r)/2$ as the final similarity threshold $\delta$ (line-11); (vi) finally, use the value of $\delta$ to return the final graph subset $\mathcal{G}_{subset}$ (line-13).
Compress Function. The workflow of the $Compress()$ is shown in Appendix C and explained as follows: (i) start with an empty graph $(\mathcal{G}^{\prime})$ (line-1); (ii) iterate the information elements in the graph $(g_{i}\in\mathcal{G})$ to compute the similarity score of $g_{i}$ with the elements in $\mathcal{G}^{\prime}$ to look for maximal similarity, i.e., $sim_{max}$ (lines 2-3); (iii) compare $sim_{max}$ against the compression threshold $\delta$ to insert $g_{i}$ in $(\mathcal{G}^{\prime})$ (lines 4-5); (iv) finally, return $(\mathcal{G}^{\prime})$ as the final subset of the graph. The end-goal of Algorithm 2 is to select highly dis-similar information elements (Ali et al., 2019) by neglecting cases with $sim_{max}>\delta$. For such cases, we assume that the corresponding information element, i.e., $g_{i}=(e_{i},r_{i},e_{i}^{{}^{\prime}})$ is redundant because there is already an element in $\mathcal{G}^{\prime}$ that is very similar to $g_{i}$.
5 Experiments
In this section, we conduct comprehensive experiments for the performance evaluation for Prompt-SAW compared against different baseline models.
5.1 Experiment Settings
Datasets. To comprehensively evaluate the effectiveness of compressed prompts, we evaluate their performance on task-agnostic and task-aware datasets. For task-agnostic dataset, we propose GSM8k-aug, i.e., an extended variant of the original GSM8K (Cobbe et al., 2021) devised by us to report the model performance under $i$-shot settings with $i\in\{1,2,4\}$ (see Appendix D.1 for details). Note, GSM8k-aug has a broader coverage, as it encompasses the experimental settings of the current GSM8K data settings (i.e., 1-shot). For the task-aware dataset, we use NaturalQuestions (Liu et al., 2023). The statistics of dataset is given in Table 5, and their detailed description is provided in Appendix D.1.
Baselines. We compare the performance of Prompt-SAW against following models as baselines: (i) Selective-Context (Li, 2023), (ii) LLMLingua (Jiang et al., 2023a), (iii) LongLLMlingua (Jiang et al., 2023b), and (iii) GPT4 (Achiam et al., 2023). Details about the baselines are provided in Appendix D.2.
Evaluation Metrics. For GSM8k-aug, we use Exact Match (EM) as the evaluation metric. This metric is also employed by (Cobbe et al., 2021) and (Jiang et al., 2023a). For the evaluation of NaturalQuestions data, we used Span Accuracy (Span-Acc) as a metric. This is similar to previous work by (Liu et al., 2023) and (Jiang et al., 2023b). Apart from this, we also use fluency (FL) (Meng et al., 2022) to measure the readability and grammatical coherence of the compressed prompt. The details about these metrics are given in Appendix D.3.
(GSM8k-aug)
(1-shot) (2-shot) (4-shot)
Method EM Tokens $1/\eta$ EM Tokens $1/\eta$ EM Tokens $1/\eta$
Original No.1 (1-shot) 71.04 422 1.00
No.2 (1-shot) 74.98 311 1.00 76.81 733 1.00
No.3 (1-shot) 73.92 249 1.00
No.4 (1-shot) 73.39 242 1.00 79.53 491 1.00 78.92 1224 1.00
Selective-Context No.1 (1-shot) 53.28 317 1.33
No.2 (1-shot) 54.74 231 1.35 57.34 550 1.33
No.3 (1-shot) 58.40 190 1.32
No.4 (1-shot) 55.78 175 1.39 52.17 349 1.41 58.13 881 1.39
LLMLLingua No.1 (1-shot) 68.43(0.5%) 321 1.32
No.2 (1-shot) 60.02 234 1.33 65.71 498 1.47
No.3 (1-shot) 66.53 202 1.23
No.4 (1-shot) 62.38(4.7%) 172 1.41 59.29 358 1.37 63.41(13.7%) 906 1.35
GPT4-Generation No.1 (1-shot) 62.86 283 1.49
No.2 (1-shot) 65.71(9.5%) 221 1.41 68.57(4.3%) 469 1.56
No.3 (1-shot) 66.57(1.1%) 199 1.25
No.4 (1-shot) 60.18 159 1.52 67.00(8.2%) 354 1.39 62.68 956 1.28
Prompt-SAW(Ours) No.1 (1-shot) 68.78 292 1.45
No.2 (1-shot) 71.98 212 1.47 71.53 469 1.56
No.3 (1-shot) 67.27 160 1.56
No.4 (1-shot) 65.32 165 1.47 72.47 329 1.49 72.12 820 1.49
Table 1: Experimental results on GSM8k-aug. We report the number of tokens in original and compressed prompts along with EM and compression ratio $(1/\eta)$. We bold-face the overall best-performing scores with existing state-of-art underlined.
Experimental Setup. In this work, we select GPT-3.5-Turbo-0613 as the target LLM, which is accessible via OpenAI API.  [^1] https://openai.com/blog/openai-api Following the setting of LLMLingua (Jiang et al., 2023a), we employ greedy decoding with the temperature set to 0. For graph construction, we use open-source language model  Vicunna-7B (Chiang et al., 2023). We use OpenAI embedding API  [^2] https://openai.com/ as embedding encoder $(E)$. The value for $\eta^{*}$ is set to {0.1, 0.2, 0.3, 0.4, 0.5} for both  GSM8k-aug and NaturalQuestions. Value of $\gamma$ in Algorithm 2 is 0.001. All the results reported in this paper are averaged over five runs. All experiments were performed using PyTorch 2.1.0 with Nvidia RTX 4090 24GB GPU.
5.2 Experimental Results
Results for Task-agnostic Settings. We first study the performance of Prompt-SAW for task-agnostic settings, whose results on GSM8k-aug are shown in Table 1. Unlike existing research that reports their performance for the 1-shot setting, we report these results for 1-shot, 2-shot and 4-shot settings, where $i$-shot indicates $i$ numbers of prompts have been employed by Prompt-SAW.
Comparing these results against the baseline models, we can observe that Prompt-SAW outperforms the previous state-of-the-art by a significant margin. For instance, compared to the best performing baselines, Prompt-SAW improves the EM score by up to 9.5%, 8.2%, and 13.7% under 1-shot, 2-shot and 4-shot settings, respectively. Correspondingly reduction in the prompt size is 31.8%, 33.0%, and 33.0%. We attribute such drastic performance improvement to the following factors: (1) Prompt-SAW retains the logical integrity of the COT prompts by sub-dividing the original prompts into smaller comprehensive information elements; (2)Prompt-SAW benefits from the workflow that allows selecting and omitting individual information elements for prompt compression without destroying the overall information structure of the compressed prompt. These help Prompt-SAW to ensure the utility of the compressed prompt for the end task.
Results for Task-aware Settings. Table 2 reports the performance of Prompt-SAW under task-aware settings on NaturalQuestions. These results show that, Prompt-SAW improves the Span Accuracy by 14.30%, 10.62% and 6.52%, respectively, for different values of the target compression rates, i.e., $\eta^{*}=\{0.5,0.3,0.1\}$, against the best-performing baseline (GPT-4). Correspondingly, the reduction in the prompt size is 56.7%, 74.0%, and 93.7%.
Correlating the results for both settings, we observe that compared to the task-agnostic scenarios, Prompt-SAW yields much stable performance for the task-aware settings. This is due to the fact that it is more difficult to dig out the latent correlation between information elements within the prompt’s internal structure rather than explicit task-aware correlation extraction. From Table 2 and Figure 3 we can also find that the performance of Prompt-SAW drops when the value for the target compression rate $(\eta^{*})$ decreases from 0.5 to 0.1. A probable justification that the actual compression ratio of Prompt-SAW is significantly higher than the target compression ratio when the target compression ratio is 10 (i.e., $\eta^{*}=0.1$). This is owing to the fact that Prompt-SAW only retains the information elements in $\mathcal{G}$ as the key/basic information units for the compression process. It will delete some entities and relations that may be highly similar to the problem but their overall structure is too long, leading to relatively poor performance for higher compression ratio.
Method NaturalQuestions
$\eta^{*}$ = 0.5 $\eta^{*}$ = 0.3 $\eta^{*}$ = 0.1
Span-Acc Tokens 1/$\eta$ Span-Acc Tokens 1/$\eta$ Span-Acc Tokens 1/$\eta$
Original 92.18 524 1.00 92.18 524 1.00 92.18 524 1.00
Selective-Context 49.23 283 1.85 45.71 173 3.03 31.12 68 7.69
LongLLMlingua 59.65 270 1.94 52.02 161 3.25 47.14 57 9.21
GPT4-Generation 72.55(14.30%) 252 2.08 66.19(10.62%) 121 4.35 50.76(6.52%) 84 6.25
Prompt-SAW(Ours) 82.93 227 2.31 73.22 136 3.86 54.07 33 16.08
Table 2: Experimenal results on NaturalQuestions.
Target LLMs GSM8K-AUG Natural Questions
1shot 2shot 4shot
GPT3.5 67.83 72.00 72.12 73.22
GPT4 91.48 93.17 95.92 74.36
Table 3: The performance comparison of Prompt-SAW using different target LLMs.
5.3 Further Analysis
In this section, we perform an in-depth analysis of the performance of Prompt-SAW.
Different Target LLMs. We also analyze the performance of Prompt-SAW using different target LLMs, with results reported in Table 3. We can find that GPT4 performs better than GPT3.5 under the task-agnostic setting, while they perform similarly under the task-aware setting. We attribute this result to GPT4’s stronger inference capabilities, so it can benefit more and learn how to perform mathematical reasoning from compressed prompts in a task-agnostic setting. However, when we consider the task-aware settings, the performance of GPT4 and GPT3.5 is similar. That is because the LLMs need to extract key information from the prompt to answer the question. If we compress the important information, LLM will make mistakes no matter how strong the reasoning ability it has.
Refer to caption Refer to caption
Figure 3: Different target compress rate between LLMLingua, GPT4 and Prompt-SAW.
Interpretability of Compressed Prompts. As explained in the introduction (also highlighted in Figure 1), a key limitation of existing prompt compression approaches is the limited interpretability of the compressed prompt. In order to validate the results of Prompt-SAW in terms of human readability and/or interpretability, we report some example prompts along with prompt compressed using Prompt-SAW and LLMLingua (Jiang et al., 2023a) in Appendix E.1, for a quick comparison. These examples clearly indicate that the prompt compressed by Prompt-SAW exhibit better readability and/or interpretability compared to compressed using LLMLingua. This loss of interpretability may significantly impact the end-utility of the compressed prompt.
We also conducted a quantitative comparison to support our claims. Specifically, we assess the fluency of the compressed prompts through the computation of a weighted mean of bi-gram and tri-gram entropies (Meng et al., 2022). Corresponding results in Table 4 show that Prompt-SAW yields relatively higher fluency scores than the baseline models. A lower score for baseline models, e.g., LLMLingua, may be attributable to loss of intrinsic semantic relationship between the tokens for the compressed prompt.
Selective-Context LLMLingua GPT4 Prompt-SAW
FL 5.61 5.74 6.07 6.30
Table 4: Fluency(FL) of the compressed prompt on GSM8k-aug. We report the performance of Prompt-SAW compared against the baseline models. The best scores are bold-faced.
6 Conclusion
In this work, we propose Prompt-SAW that leverages graph structures to infer key information in the prompt in order to come up with a compressed prompt. Experimental evaluation shows that Prompt-SAW outperforms the existing research on prompt compression by a significant margin. Moreover, Prompt-SAW also addressed a key limitation of existing prompt compression approaches, i.e., the compressed prompts are human-friendly and easy to understand. In the future, we also plan to extend this work to other languages by leveraging techniques for mapping embedding vectors (Ali et al., 2023a; b).
Appendix A Background
A.1 Knowledge Graph
Knowledge Graph (KG) can be represented as $\mathcal{G}=\{(s,r,o)\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations. Here entities are represented as nodes in $\mathcal{G}$, while relations $\mathcal{R}$ form up edges between the nodes.
A.2 Task-aware Prompts
Task-aware prompts refer to the ones that are strongly related to the task and need to be re-compressed while changing the question. These prompts usually contain the specific information needed to solve the task, and some redundant parts can be removed. For example, the task can be the question, ”Who are the first people to win the Nobel Prize?” and the prompt may contain a document that includes all the information about people who won the Nobel Prize.
A.3 Task-agnostic Prompts
Opposite to task-aware prompts, task-agnostic prompts are weakly related to the task. This kind of prompt usually just provides LLMs with an example of what to do, such as how to solve the problem step by step. For example, the prompt may be ”The weather is really nice today, emotion: positive.” The model then follows this format and judges the emotion of the input sentence.
A.4 Chain-of-thought Prompt
This is a type of task-agnostic prompt. Chain-of-thought Prompt aims to improve performance on tasks requiring logic and calculation by mimicking human reasoning. That is $x^{info}$ include several demonstrations with detailed reasoning $p^{info}=\{p^{demo}_{1},p^{demo}_{2},\cdots\}$. We research how to compress chain-of-thought prompts on mathematical reasoning tasks.
A.5 Prompt with external documentation
This is a type of task-aware prompt. Prompt with external documentation implies that the prompt contains some additional information, such as external knowledge, that the model may refer to to answer the question. That external knowledge $x^{info}$ may include several external documents $p^{info}=\{p^{doc}_{1},p^{doc}_{2},\cdots\}$. Question Answering is a specific scenario for prompts with external knowledge.
A.6 Token-level Prompt compression
In this section, we introduce the previous token-level compression method, e.g., LLMlingua (Jiang et al., 2023a), and LongLLMlingua (Jiang et al., 2023b).
LLMLingua is a token-level prompt compression method that performs compression based on perplexity. It includes a budget controller to calculate the compression ratio of demonstrations and the further compression ratio of each demonstration based on given parameters. LLMLingua uses the LLAMA-2 (Touvron et al., 2023) to calculate the perplexity of each token. Finally, LLMLingua compresses the prompt based on the perplexity and compression ratio.
LongLLMlingua is a token-level prompt compression method that aims at task-aware prompt compression. It changes the perplexity measure method and relates it to the specific question.
Appendix B Prompts
B.1 Prompts for Graph Construction
Prompts for Graph Construction: Example:
Input:
Deadpool 2 is scheduled to be released in the United States on May 18, 2018. A sequel, Deadpool 3, is in development.
Output:
Deadpool 2; is scheduled to be released in; the United States on May 18, 2018$>$
$<$Deadpool 3; is in; development$>$ Hint: You should only respond the knowledge graph triplet and not contain other word. The knowledge graph triplet is formulated as $<s,r,o>$, $s$ and $o$ should not be too long. Please keep all the relations atomic and indivisible. Please generate the entity and relation triplets of the Input: Input:
B.2 Instructions used for GPT-4 response Generation
The instructions we used in the GPT-4 Generation are shown below:
Instructions used for GPT-4 response Generation: Instruction1. Condense the given paragraph to just 50% of its original size, focusing on the core message. Instruction2. Reduce the length of the specified paragraph to 50%, keeping only the most essential information. Instruction3. Compress the paragraph to 50% of its length, ensuring the main idea is intact. Let’s do it step by step. Instruction4. You are a prompt compression expert. Please compress the following prompt to 50% of its original length. Let’s do it step by step. Instruction5. You are a prompt compression expert. Please compress the following prompt with the following steps: (1) Find the key information of the document (2) Compress the prompt to 50% of its original length without damaging key information. Let’s do it step by step.
Appendix C Additional Implementation details
Algorithm 3 Compress Task-agnostic Prompt
1:
2: #$\delta:$ Compression threshold
[…]
4: #$E:$ encoder
5: #$sim():$ function used to calculate similarity
6:subgraph $\mathcal{G}^{{}^{\prime}}$
7:$\mathcal{G}^{\prime}=\{\}$
8:for $g_{i}\in\mathcal{G}$ do
9:     $sim_{max}=max\{sim(E(g),E(g_{i}))\;\;\forall g\in\mathcal{G}^{\prime}\}$
10:     if $sim_{max}<=\delta$ then
11:         $\mathcal{G^{\prime}}.insert(g_{i})$
12:     end if
13:end for
14:return $\mathcal{G}^{{}^{\prime}}$
Appendix D Experimental Details
D.1 Dataset
We provide a detailed description of the evaluation data sets below. The statistics of the dataset are given in Table 5.
statistics GSM8K-AUG NaturalQuestions
No.1 shot No.2 shot No.3 shot No.4 shot
Token number of prompt 422 311 249 242 3040
Number of questions 1319 1319 1319 1319 2654
Table 5: The statistics of the dataset
(i) GSM8k-aug. Originally GSM8K (Cobbe et al., 2021) data set encompasses 8,500 high-quality, linguistically diverse grade school math word problems. We find that existing research poses some limitations on prompt compression experiments GSM8k, i.e., they only allow compressing one COT prompt (1-shot). On the contrary, we design a more comprehensive experimental setting by extending it to $i$-shot setting ($i\in\{1,2,4\}$), i-shot means $i$ chain-of-thought demonstrations in the prompt. However, this setting may result in a huge amount of experiments. To avoid computational overhead, we only choose a small subset of questions in GSM8K for testing (i.e., approximately 30%). In this paper, we call this setting as GSM8k-aug.
(ii) NaturalQuestions. It is a QA dataset that is comprised of real-world queries collected by individuals (Liu et al., 2023). Each question of this dataset has 20 related documents, one of which contains the correct answer. We select documents containing answers as compression targets to examine better the compression performance of different methods on a single document.
D.2 Baseline Models
(i) Selective-Context. Selective-Context by Li (2023) uses a small language model to calculate the self-information in the prompt and then filter out on token-level based on the self-information of each token.
(ii) LLMLingua. LLMLingua by Jiang et al. (2023a) perform token-level prompt compression based on the perplexity calculated by the small language model.
(iii) LongLLMLingua. Based on LLMLingua, LongLLMlingua by Jiang et al. (2023b) further adds a coarse-grained filtering module, which is more suitable for long document compression.
We followed their original experimental setting, uses LLMLlingua on GSM8K and LongLLMlingua on NaturalQuestions.
(iii) GPT4.(Achiam et al., 2023))
We designed five sets of prompts for GPT4 to inspire its ability on prompt compression and reported the best one. Appendix B displays the prompts we employed.
D.3 Evaluation Metrics
Detailed description and mathematical formulation of the evaluation metrics is provided as follows:
Exact Match (EM). In EM, when the model output answer is completely consistent with the golden answer, the answer is considered correct. It is shown below.
$\mathbbm{1}\left[\bigvee_{q\in\mathcal{Q}}[f(compress(q))=q^{*}]\right]$ (6)
Where $f(\cdot)$ represents the model used to answer the question, $\mathcal{Q}$ and $q^{*}$ represent the question, and $q^{*}$ indicate the answer for question, and $compress$ indicate the prompt compression method.
Span Accuracy (SAcc). We follow previous work and use SAcc to measure the performance of QA datasets. SAcc determines whether the standard answer is part of the response answer of the GPT model, as shown below.
$\mathbbm{1}\left[\bigvee_{q\in\mathcal{Q}}[q^{*}\in f(compress(q))]\right]$ (7)
Fluency (FL). We use fluency as a metric to measure the readability and grammatical coherence of the compressed prompt. Following Meng et al. (2022), we use the following formula to compute the fluency.
$\text{FL}=-\sum_{k}f(k)\log_{2}f(k)$ (8)
where ${f(\cdot)}$ means the $n$-gram frequency distribution.
Appendix E Additional Results
E.1 Interpretability of Compressed prompts (Examples)
In this section, we report some examples prompts along with prompts compressed by Prompt-SAW and LLMLingua (Jiang et al., 2023a) for a quick comparison in terms of readability and/or interpretability of the compressed prompt. As an example, for the compressed prompt text compressed using LLMLingua in Table E.1, the text ”won twoes forg01 theate women prize:ertMayer” is hard to interpret for humans. On the contrary, the prompt compressed by Prompt-SAW yields comprehensive information units helpful that are not only easy to interpret but are also highly relevant to the question.
Example 1. Original Prompt:
Write a high-quality answer for the given question using only the provided search results.
Document [1](Title: List of Nobel laureates in Physics) The first Nobel Prize in Physics was awarded in 1901 to Wilhelm Conrad Röntgen, of Germany, who received 150,782 SEK, which is equal to 7,731,004 SEK in December 2007. John Bardeen is the only laureate to win the prize twice—in 1956 and 1972. Maria Skłodowska-Curie also won two Nobel Prizes, for physics in 1903 and chemistry in 1911. William Lawrence Bragg was, until October 2014, the youngest ever Nobel laureate; he won the prize in 1915 at the age of 25. Two women have won the prize: Curie and Maria Goeppert-Mayer (1963). As of 2017, the prize has been awarded Question: who got the first nobel prize in physics. Answer: Compressed Prompt by LLMLingua:
Write a high-quality answer for the given question using only the provided search results.
1Title: List of Nobelates in The first Prize1 Wilhelmrad, of who received82 in en prize. won twoes forg01 theate women prize:ertMayer (1963). As of 2017, the prize has been awarded Question: who got the first nobel prize in physics. Answer: Compressed Prompt by Prompt-SAW:
Write a high-quality answer for the given question using only the provided search results.
Wilhelm Conrad Röntgen awarded first Nobel Prize in Physics 1901.William Lawrence Bragg won Nobel Prize in Physics 1915.Maria Goeppert-Mayer won Nobel Prize in Physics 1963 Question: who got the first nobel prize in physics. Answer:
Example 2. Original Prompt:
Write a high-quality answer for the given question using only the provided search results.
Document [1](Title: Distilled beverage) The term s̈piritïn reference to alcohol stems from Middle Eastern alchemy. These alchemists were more concerned with medical elixirs than with transmuting lead into gold. The vapor given off and collected during an alchemical process (as with distillation of alcohol) was called a spirit of the original material. Question: where did the term spirits for alcohol come from Answer: Compressed Prompt by LLMLingua:
Write a high-quality answer for the given question using only the provided search results.
) was called a spirit of the original material. Question: where did the term spirits for alcohol come from Answer: Compressed Prompt by Prompt-SAW:
Write a high-quality answer for the given question using only the provided search results .
Alchemical process involves distillation of alcohol.Spirit stems from Middle Eastern alchemy Question: where did the term spirits for alcohol come from Answer:
Example 3. Original Prompt:
Write a high-quality answer for the given question using only the provided search results.
Document [1](Title: OPEC) Organization of the Petroleum Exporting Countries (OPEC, /oupek/ OH-pek, or OPEP in several other languages) is an intergovernmental organization of 14 nations as of February 2018, founded in 1960 in Baghdad by the first five members (Iran, Iraq, Kuwait, Saudi Arabia, and Venezuela), and headquartered since 1965 in Vienna, Austria. As of 2016, the 14 countries accounted for an estimated 44 percent of global oil production and 73 percent of the world’s p̈rovenöil reserves, giving OPEC a major influence on global oil prices that were previously determined by American-dominated multinational oil companies. Question: how many countries are a part of opec Answer: Compressed Prompt by LLMLingua:
Write a high-quality answer for the given question using only the provided search results.
Title: OPE ofC, /0̆2c8kkPEP in otheral1 nations as1 in by Venezuela. of the4 on by Americanatedinational oil companies. Question: how many countries are a part of opec Answer: Compressed Prompt by Prompt-SAW:
Write a high-quality answer for the given question using only the provided search results.
Organization of the Petroleum Exporting Countries abbreviation OPEC.Organization of the Petroleum Exporting Countries nations involved 14.Organization of the Petroleum Exporting Countries founded in 1960 Question: how many countries are a part of opec Answer:
Example 4. Original Prompt:
Write a high-quality answer for the given question using only the provided search results.
Document [1](Title: Subcutaneous injection) A subcutaneous injection is administered as a bolus into the subcutis, the layer of skin directly below the dermis and epidermis, collectively referred to as the cutis. Subcutaneous injections are highly effective in administering vaccines and medications such as insulin, morphine, diacetylmorphine and goserelin. Subcutaneous, as opposed to intravenous, injection of recreational drugs is referred to as s̈kin popping.̈ Subcutaneous administration may be abbreviated as SC, SQ, sub-cu, sub-Q, SubQ, or subcut. Subcut is the preferred abbreviation for patient safety. Question: where would a subcutaneous injection be made in the skin Answer: Compressed Prompt by LLMLingua:
Write a high-quality answer for the given question using only the provided search results.
Document [1](Title: Subcutaneous injection) A subcutaneous injection is administered as a bolus into the subcutis, the layer of skin directly below the dermis and epidermis, collectively referred to as the cutis. Subcutaneous injections are highly effective in administering vaccines and medications such as insulin, morphine, diacetylmorphine and goserelin. Subcutaneous, as opposed to intravenous, injection of recreational drugs is referred to as s̈kin popping.̈ Subcut SubQ, or subcut. Subcut is the preferred abbreviation for patient safety Question: where would a subcutaneous injection be made in the skin Answer: Compressed Prompt by Prompt-SAW:
Write a high-quality answer for the given question using only the provided search results.
Subcutaneous injection administered as bolus into the subcutis.Subcutaneous injection administered for vaccines and medications.Subcutaneous injection referred to as s̈kin popping\̈\ Question: where would a subcutaneous injection be made in the skin Answer:
Appendix F Limitations
Although Large Language Models (LLMs) have garnered widespread attention for their remarkable capacity for knowledge comprehension (Yang et al., 2024b; a), enabling tailored solutions across various applications, they also face critical issues such as privacy concerns (Hu et al., 2023a), and explainability (Hu et al., 2023b; Lai et al., 2023). LLM applications often involve data containing sensitive information, necessitating effective solutions to safeguard privacy (Xu et al., 2023a). One promising approach to address this challenge is through the design of Differentially Private (DP) algorithms (Dwork et al., 2006). DP offers provable protection against identification and is resilient to arbitrary auxiliary information that may be available to attackers. While there have been numerous studies on DP machine learning (Hu et al., 2022; Wang et al., 2020; 2021; Su et al., 2022; Hu et al., 2023c; Wang et al., 2023) and DP deep learning (Xiang et al., 2024; 2023; Shen et al., 2023), most of these efforts have primarily focused on either continuous tabular data or image data. Unfortunately, there has been less emphasis on adapting DP algorithms to the context of Natural Language Processing (NLP) and the text domain. Addressing this gap is paramount as text data presents unique challenges and characteristics that necessitate specialized privacy-preserving techniques. By developing and refining DP algorithms tailored to NLP tasks, we can enhance the privacy protections of LLMs and facilitate their responsible and ethical deployment across various domains. However, this endeavor is left for future exploration and research.